# FDGaussian: Fast Gaussian Splatting from Single Image via Geometric-aware Diffusion Model

### v1

@@ 3.4 Epipolar Attention for Multi-view Rendering
Consider the intermediate UNet feature $f_{s}$, we can compute its corresponding epipolar lines $\{l_{t}\}_{t\neq s}$ on the feature map of all other views $\{f_{t}\}_{t\neq s}$ (please refer to Supplementary Materials for the details). Each point $p$ on $f_{s}$ will only access the features that lie along the camera ray (in other views) as all points in its own views during rendering. We then estimate the weight maps for all positions in $f_{s}$, stack these maps, and get the epipolar weight matrix $M_{st}$. Finally, the output of the epipolar attention layer $\hat{f}_{s}$ can be formulated as:
@@
In this way, our proposed epipolar attention mechanism facilitates the efficient and accurate association of features across multiple views. By constraining the search space to the epipolar lines, we effectively reduce the computation cost as well as eliminate potential artifacts.
3.5 Loss Function
During the training of the reconstruction stage, we suppose each reference image $\mathcal{I}_{ref}$ has $N$ corresponding views $\mathcal{I}$ with the relative camera pose change $\Delta\pi$. Then we feed the reference image $\mathcal{I}_{ref}$ into the network, and minimize the average reconstruction loss of target view $\mathcal{I}^{(s)}$:
$\mathcal{L}_{rec}=\frac{1}{N}\sum_{s=1}^{N}\|\mathcal{I}^{(s)}-g(f(\mathcal{I}_{ref}),\Delta\pi^{(s)})\|^{2},$ (7)
@@ 4 Experiments
4.1 Experimental Setup
4.1.1 Implementation Details
The two stages are trained separately. For the generation stage, we use the similar network structure of [30, 32] and initialize the weight with Zero-1-to-3 pre-trained weights for training efficiency. We utilize a Vision Transformer (ViT) model of depth 6 as the reference image encoder and generate an output of size $1024\times 256$. The decoding process involves two decoders, i.e. image plane decoder and orthogonal plane decoder, each with a depth of three and outputs a feature map $F\in\mathbb{R}^{128\times 128\times 64}$. After the multi-view generation, we directly adopt the implementation of [67] to select 16 views with the highest perceptual quality score. For the reconstruction stage, the network that maps the input images to the mixtures of Gaussians is architecturally identical to the UNet[46]. The last layer is replaced with a $1\times 1$ convolutional layer with 15 output channels. As mentioned in Sec. 3.4, in order to allow the network to coordinate and exchange information between views, we add epipolar attention blocks after residual blocks followed by the cross-attention layers. We use the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ with a learning rate of $10^{-4}$. All experiments are performed and measured under NVIDIA V100 (16GB) GPU.
Figure 5: Qualitative comparison of 3D reconstruction results with baselines.
4.1.2 Datasets
@@
We train our diffusion model on the recently released Objaverse[11] dataset, which is a large-scale CAD dataset containing 800K high-quality objects. We directly employ the processed rendering data from Zero-1-to-3, which provides 12 random views of each object. For evaluation, we use the test split of Objaverse provided by Zero-1-to-3. In addition, to test the performance of our model on the out-of-distribution data, we also evaluate the Google Scanned Object dataset[12], which contains high-quality scanned household items. During the training stage, images are resized to $256\times 256$ resolution.
4.1.3 Baselines
We mainly evaluate our approach against methods that can generalize to open-set categories and accept single-view RGB images as inputs. In particular, we adopt Zero-1-to-3[30], Realfusion[34], Consistent-123[68], Shap-E[20], and DreamGaussian[50] as baseline methods. Zero-1-to-3 is able to synthesize novel views conditioned on viewpoints without training data. Realfusion is based on Stable Diffusion and utilizes SDS loss for single-view reconstruction. Shap-E converts the single input image into a point cloud encoded in MLP and is trained on the OpenAI 3D dataset. DreamGaussain leverages 3D Gaussian Splatting and diffusion priors during reconstruction, which greatly improves the speed. We adopt the implementation of ThreeStudio[17] for reconstruction with Zero-1-to-3, which achieves better performance than the original implementation. For other works, we use their officially released code for quantitative and qualitative evaluation.
4.1.4 Evaluation metrics
@@
We mainly focus on two tasks, i.e. novel view synthesis (NVS) and single image 3D reconstruction (1-to-3). On the NVS task, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM)[54], and Learned Perceptual Image Patch Similarity (LPIPS)[71] to measure the similarity between rendered images and ground truth images. On the 1-to-3 task, we report the commonly used Chamfer Distance (CD) and CLIP similarity[41].
4.2 Main Results
4.2.1 Comparison on novel view synthesis
@@
As shown in Tab. 1, FDGaussian surpasses all baseline methods regarding PSNR, LPIPS, and SSIM, indicating it provides a sharper and more accurate reconstruction. The qualitative result of our method is demonstrated in Fig. 3. The nearby views synthesized by FDGaussian are geometrically and semantically similar to the reference view, while the views with large viewpoint change showcase reasonable diversity. Furthermore, the orthogonal-plane decomposition mechanism enables our model to capture the details of the input image (discussed in Sec. 4.3).
Table 1: The quantitative comparison for novel-view synthesis. We color each cell as best and second best. We report PSNR, SSIM, and LPIPS on Objaverse[11] and GSO[12] datasets. The proposed FDGaussian significantly improves the view consistency compared with baselines by a large margin.
Objaverse Google Scanned Objects
@@
Zero-1-to-3[30] 18.68 0.883 0.189 18.37 0.877 0.212
Realfusion[34] 18.95 0.882 0.167 15.26 0.722 0.283
Consistent-123[68] 20.72 0.877 0.122 19.46 0.858 0.146
DreamGaussian[50] 21.53 0.915 0.122 19.93 0.895 0.177
FDGaussian(Ours) 23.97 0.921 0.113 22.98 0.899 0.146
Table 2: The quantitative comparison for single-view 3D reconstruction. We report CLIP similarity, Chamfer Distance and the (overall) reconstruction time on GSO[12] dataset.
@@
Shap-E[20] 68.4 0.0636 1min
Zero-1-to-3[30] 79.1 0.0339 30min
Realfusion[34] NeRF-based 71.5 0.0819 20min
DreamGaussian[50] 75.8 0.0246 2min
FDGaussian(Ours) GS-based 80.0 0.0232 70s
4.2.2 Comparison on 3D reconstruction
For the single-image 3D reconstruction task, we show the results in Tab. 2. FDGaussian outperforms competing approaches by a substantial margin. By leveraging the pruning techniques, we further reduce the overall reconstruction time to about 70 seconds. Fig. 5 displays the qualitative comparison results between our method and the baselines. From the visual comparison, we discover that Shap-E is susceptible to collapsing during the generation process, resulting in an unpredictable and blurring outcome. Zero-1-to-3 suffers from multi-view inconsistency. Although DreamGaussian keeps the generated results consistent and semantically faithful to the reference image, it fails to synthesize intricate details and often produces over-smooth textures. This demonstrates the superiority of FDGaussian over the current state-of-the-art methods and its capacity to generate high-quality 3D objects even with complex structures.
Figure 6: Qualitative ablation study of different components.
@@ 4.3 Ablations and Analyses
4.3.1 Overall ablation study
FDGaussian is composed of two stages: geometric-aware multi-view generation and the following Gaussian Splatting reconstruction. We present the qualitative ablation results in Fig. 6. The orthogonal-plane decomposition mechanism plays a crucial role in generating geometrically consistent novel views. CLIP embedding also helps to eliminate the artifacts of the synthesized views while preserving the semantic fidelity to the reference image. The epipolar attention demonstrates great importance in fusing images from various viewpoints. Without it, even highly consistent multi-view images can lead to inconsistent 3D results.
4.3.2 Ablations of multi-view generation
Our multi-view generation stage mainly consists of geometric and semantic guidance. Removing them respectively or simultaneously gives us four different combinations. As shown in Tab. 3 and Fig. 6, the orthogonal-plane decomposition mechanism contributes the most to the geometric accuracy and consistency, bringing about visual enhancement to a great extent. The semantic guidance further increases the metric score and slightly improves visual consistency.
Table 3: Ablation studies of multi-view generation. Evaluated on the GSO dataset[12].
@@ 4.3.2 Ablations of multi-view generation
4.3.3 Number of synthesized views
We adopt the Chamfer distance (CD) to evaluate the quality of reconstruction. As shown in Tab. 5, we find that as the number of synthesized views increased, the quality of 3D reconstruction is enhanced accordingly but the time required does not increase much. This is expected since there is more overlapping and interaction across views. However, when the number of views reaches 32, the total time increases dramatically while the quality improvement is insignificant. This might be because over-sufficient views could become the bottleneck of computation.
Table 4: The quantitative comparison of different numbers of synthesized views. Here the number of views includes the reference view. The generation time refers to the time of multi-view generation.
# of views CD gene. time recon. time
4 0.0552 9s 52s
8 0.0327 10s 53s
16 0.0233 15s 55s
32 0.0232 21s 68s
Threshold CD recon. time
w/o GDS 0.0234 15min
0.01 0.0232 93s
0.1 0.0233 55s
0.5 0.0235 78s
Table 4: The quantitative comparison of different numbers of synthesized views. Here the number of views includes the reference view. The generation time refers to the time of multi-view generation.
Table 5: Quantitative results of ablating GDS metric. A proper choice of GDS threshold leads to significant efficiency improvement.
4.3.4 Acceleration of the optimization
As mentioned in Sec. 3.3, we propose to use the Gaussian Divergent Significance (GDS) measure to further regularize the split and clone process. As demonstrated in Tab. 5, this strategy has significantly reduced the optimization time while not sacrificing the reconstruction quality, leading to at most 15$\times$ faster convergence speed when compared with the original split and clone operation proposed in [21].
Figure 7: Text-to-3D. FDGaussian, when combined with text-to-image models[44, 43], enables text-to-3D.
4.4 Compatibility with Text-to-Image Models
FDGaussian seamlessly integrates with the off-the-shelf text-to-image models[44, 43]. These models convert textual descriptions into 2D images, which our model further transforms into high-quality multi-view images and Gaussian representations. Visual examples are shown in Fig. 7. Notably, our model excels in reconstructing the essence of the given 2D image, even managing to capture details for occluded parts.
@@ 4.5 Limitations and Future Works
While FDGaussian shows promising results in reconstructing 3D objects from single-view images, there are still some limitations that the current framework does not entirely address. First, the number of generated views is fixed in our method. Adaptively generating different numbers of views for objects with different topological symmetries might further reduce the total reconstruction time. Additionally, our current method is restricted to single-object 3D reconstruction. It remains to be extended to complex scenes or multi-object reconstruction in the future.
@@ 5 Conclusions
In this work, we proposed a two-stage model named FDGaussian to reconstruct 3D objects from single-view images. This method first synthesizes consistent yet 3D-aware multi-view images via a diffusion model under the guidance of an orthogonal-plane decomposition mechanism. Then epipolar attention is leveraged to render with these images during Gaussian Splatting. The novel metric, i.e. Gaussian Divergent Significance (GDS), is proposed to accelerate optimization. Qualitative and quantitative results show that the proposed method reconstructs 3D Gaussian representations that 1) are consistent in different viewpoints, 2) are high fidelity to the reference image, and 3) display plausible creativity in the unseen areas.
Appendix 0.A Appendix
0.A.1 Details of Gaussian Divergent Significance (GDS)
@@
$\Upsilon_{GDS}(G(\boldsymbol{x}_{1}),G(\boldsymbol{x}_{2}))=\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|^{2}+tr(\boldsymbol{\Sigma}_{1}+\boldsymbol{\Sigma}_{2}-2(\boldsymbol{\Sigma}_{1}^{-1}\boldsymbol{\Sigma}_{2}\boldsymbol{\Sigma}_{1}^{-1})^{1/2}).$ (9)
The first term $\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|^{2}$ can be expressed as $(\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2})^{T}(\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2})$. Additionally, the covariance matrix can be factorized into a scaling matrix $S$ and a rotation matrix $R$: $\boldsymbol{\Sigma}=RSS^{T}R^{T}$. Since the scaling matrix is a diagonal matrix and the rotation matrix is orthogonal, we can thus simplify the computation of the second term. Given:
$\boldsymbol{\Sigma}_{1}=R_{1}S_{1}S_{1}^{T}R_{1}^{T}$ (10)
$\boldsymbol{\Sigma}_{2}=R_{2}S_{2}S_{2}^{T}R_{2}^{T},$
@@
$\quad\;tr(\boldsymbol{\Sigma}_{1}+\boldsymbol{\Sigma}_{2}-2(\boldsymbol{\Sigma}_{1}^{-1}\boldsymbol{\Sigma}_{2}\boldsymbol{\Sigma}_{1}^{-1})^{1/2})$ (11)
$=tr(\boldsymbol{\Sigma}_{1})+tr(\boldsymbol{\Sigma}_{2})-tr(2(\boldsymbol{\Sigma}_{1}^{-1}\boldsymbol{\Sigma}_{2}\boldsymbol{\Sigma}_{1}^{-1})^{1/2})$
$=tr(R_{1}S_{1}S_{1}^{T}R_{1}^{T})+tr(R_{2}S_{2}S_{2}^{T}R_{2}^{T})-tr(2(\boldsymbol{\Sigma}_{1}^{-1}\boldsymbol{\Sigma}_{2}\boldsymbol{\Sigma}_{1}^{-1})^{1/2})$
$=tr(S_{1}S_{1}^{T})+tr(S_{2}S_{2}^{T})-tr(2(\boldsymbol{\Sigma}_{1}^{-1}\boldsymbol{\Sigma}_{2}\boldsymbol{\Sigma}_{1}^{-1})^{1/2})$
@@
0.A.2 Details of Epipolar Line
Figure 8: More qualitative results on wild images.

### v2

@@ 3.2 Geometry-aware Multi-view Generation
3.3 Reconstruction with Epipolar Attention
During the reconstruction stage, we target to exploit the synthesized consistent multi-view images for restoring high-quality 3D objects. However, relying solely on cross-attention to communicate between images of multiple viewpoints is insufficient. Therefore, we propose epipolar attention to allow association between the features of different views. The epipolar line for a given feature point in one view is the line on which the corresponding feature point in the other view must lie, based on the known geometric relationship between two views. It acts as a constraint to reduce the number of potential pixels in one view that can attend to another view. We present the illustration of epipolar line and epipolar attention in Figure 3.
Consider the intermediate UNet feature $f_{s}$, we can compute its corresponding epipolar lines $\{l_{t}\}_{t\neq s}$ on the feature map of all other views $\{f_{t}\}_{t\neq s}$ (please refer to the Appendix B for the details). Each point $p$ on $f_{s}$ will only access the features that lie along the camera ray (in other views) as all points in its own views during rendering. We then estimate the weight maps for all positions in $f_{s}$, stack these maps, and get the epipolar weight matrix $M_{st}$. Finally, the output of the epipolar attention layer $\hat{f}_{s}$ can be formulated as:
@@ 3.3 Reconstruction with Epipolar Attention
Our proposed epipolar attention mechanism limits the search space for corresponding features across different views, facilitating the efficient and accurate association of features across multiple views. In this way, we effectively reduce the computation cost as well as eliminate potential artifacts.
Training Objective
Given the input image $\mathcal{I}_{0}$ and multi-view images $\{\mathcal{I}_{i}\}_{i=1}^{N}$ with relative camera pose change $\{\Delta\pi_{i}\}_{i=1}^{N}$, we feed them into the network, and minimize the average reconstruction loss:
$\mathcal{L}_{rec}=\frac{1}{N}\sum_{s=1}^{N}\|\mathcal{I}_{i}-g(f(\mathcal{I}_{0}),\Delta\pi_{i})\|^{2},$ (6)
@@
3.4 Accelerating the Reconstruction
However, the split and clone operations proposed by the original work (Kerbl et al., 2023) overlook the distance between 3D Gaussians, during the optimization process which significantly slows down the process. We observe that if two Gaussians are close to each other, even if the positional gradients are larger than a threshold, they should not be split or cloned since these Gaussians are updating their positions. Empirically, splitting or cloning these Gaussians has negligible influence on the rendering quality as they are too close to each other. For this reason, we propose Gaussian Divergent Significance (GDS) as a measure of the distance of 3D Gaussians to avoid unnecessary splitting or cloning:
$\Upsilon_{GDS}(G(\bm{x}_{1}),G(\bm{x}_{2}))=\|\bm{\mu}_{1}-\bm{\mu}_{2}\|^{2}+tr(\bm{\Sigma}_{1}+\bm{\Sigma}_{2}-2(\bm{\Sigma}_{1}^{-1}\bm{\Sigma}_{2}\bm{\Sigma}_{1}^{-1})^{1/2}),$ (8)
where $\bm{\mu}_{1}$, $\bm{\Sigma}_{1}$, $\bm{\mu}_{2}$, $\bm{\Sigma}_{2}$ are the position and covariance matrix of two 3D Gaussians $G(\bm{x}_{1})$ and $G(\bm{x}_{2})$.
To be specific, we only perform the split and clone operations on the 3D Gaussians with large positional gradients and GDS. To avoid the time-consuming process of calculating GDS for every pair of 3D Gaussians, we further propose two strategies. Firstly, for each 3D Gaussian, we locate its closest 3D Gaussian by leveraging the k-nearest neighbor (k-NN) algorithm and calculate their GDS for each pair. As a result, the time complexity is reduced from $O(N^{2})$ to $O(N)$. Additionally, as mentioned in Section 3.1, the covariance matrix can be factorized into a scaling matrix $S$ and a rotation matrix $R$: $\bm{\Sigma}=RSS^{T}R^{T}$. We take advantage of the diagonal and orthogonal properties of rotation and scaling matrices to simplify the calculation of Equation 8. Details of GDS will be discussed in the Appendix A.
Figure 4: Qualitative comparisons of generated multi-view images from Objaverse dataset. The artifacts are marked with red boxes. Our method achieves better consistency and visual quality.
@@ 4 Experiments
4.1 Experimental Protocols
Datasets
@@ 4.1 Experimental Protocols
We train our multi-view generation model on the recently released Objaverse (Deitke et al., 2023) dataset, which is a large-scale CAD dataset containing 800K high-quality objects. We directly employ the processed rendering data from Zero-1-to-3, which provides 12 random views of each object. During the training, images are resized to $256\times 256$ resolution.
We evaluate the synthesized multi-view images and reconstructed 3D Gaussian Splatting (3D-GS) on the test split of Objaverse where we randomly select 100 objects beyond the training set. In addition, to test the performance of our model on the out-of-distribution data, we also evaluate the Google Scanned Object (GSO) dataset (Downs et al., 2022), which contains high-quality scanned household items.
Implementation Details
For the generation stage, we initialize the network with Zero-1-to-3 (Liu et al., 2023c) pre-trained weights for training efficiency. We utilize a Vision Transformer (ViT) model of depth 6 as the reference image encoder and generate an output of size $1024\times 256$. The decoding process involves two decoders, i.e. image plane decoder and orthogonal plane decoder, each with a depth of 3 and outputs a feature map $F\in\mathbb{R}^{128\times 128\times 64}$. After the multi-view generation, we directly adopt the implementation of Yang et al. (2022) to select 32 views with the highest perceptual quality score. For the reconstruction stage, the network that maps the input images to the mixtures of Gaussians is architecturally identical to the UNet (Song et al., 2020a). The last layer is replaced with a $1\times 1$ convolutional layer with 15 output channels. As mentioned in Section 3.3, in order to allow the network to coordinate and exchange information between views, we add epipolar attention blocks after residual blocks followed by the cross-attention layers. We use the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ with a learning rate of $10^{-4}$.
Metrics
@@
We compare generated multi-view images and rendered views from reconstructed 3D-GS with the ground truth frames, in terms of Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) (Wang et al., 2004), and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018).
Baselines
In terms of multi-view generation, we compare our GeoGS3D with Zero-1-to-3 (Liu et al., 2023c), Syncdreamer (Liu et al., 2023d), and Wonder3D (Long et al., 2024). Zero-1-to-3 finetunes Stable Diffusion to achieve novel view synthesis under a given camera pose. Syncdreamer follows a similar intuition but learns the joint distribution of all target views. Wonder3D proposes cross-domain attention to propagate information between normal and image, generating normal maps and color images simultaneously.
For image-to-3D reconstruction, we adopt DreamGaussian (Tang et al., 2023a), LGM (Tang et al., 2024a) and InstantMesh (Xu et al., 2024b). Notably, LGM and InstantMesh also adopt two-stage frameworks to achieve image-to-3D creation, which generates multi-view images and then utilizes large reconstruction models to obtain 3D objects. DreamGaussian initializes geometry and appearance using single-step SDS loss and then extracts a textured mesh with refinement guided by MSE loss.
4.2 Comparison with Existing Methods
Image to Multi-view
@@ 4.2 Comparison with Existing Methods
We first compare the generation stage of our method against recent image-to-multiview generation methods (Liu et al., 2023c; d; Long et al., 2024). The quantitative results are shown in Table 1, and the qualitative results are shown in Figure 4. GeoGS3D surpasses all baseline methods regarding PSNR, LPIPS, and SSIM, indicating it provides sharper and more accurate results. According to the qualitative results, the nearby views synthesized by GeoGS3D are geometrically and semantically similar to the reference view, while the views with large viewpoint change showcase reasonable hallucination. Furthermore, the orthogonal-plane decomposition mechanism enables our model to capture the details of the input image.
Table 1: Quantitative comparisons on the quality of generated multi-view images for image to multi-view task.
Objaverse GSO
@@
Zero-1-to-3 18.68 0.883 0.189 18.37 0.877 0.212
SyncDreamer 20.06 0.816 0.159 20.03 0.817 0.166
Wonder3D 23.52 0.895 0.130 22.69 0.883 0.147
GeoGS3D (generation) 23.97 0.921 0.113 22.98 0.899 0.146
Image to 3D
For the single-image 3D reconstruction task, we show the results in Table 2 and Figure 5. Statistically, GeoGS3D outperforms competing approaches by a substantial margin. From the visual results, our method is able to generalize to unseen data. This demonstrates the superiority of GeoGS3D over the current state-of-the-art methods and its capacity to generate high-quality 3D objects even with complex structures.
Figure 5: Qualitative comparisons for image-to-3D. The first two rows are from the Objaverse dataset, the next two are from the GSO dataset, and the final row is from an in-the-wild image. Our method demonstrates superior performance in terms of both visual fidelity and accuracy compared to existing approaches.
Table 2: Quantitative comparisons on the quality of 3D representation for image-to-3D.
Objaverse GSO
@@
DreamGaussian 19.97 0.860 0.205 20.24 0.856 0.179
LGM 20.44 0.861 0.222 18.72 0.842 0.231
InstantMesh 21.63 0.891 0.148 20.95 0.871 0.147
GeoGS3D (reconstruction) 22.35 0.894 0.143 21.75 0.896 0.141
@@ 4.3 Ablations and Analyses
The ablation studies below are conducted on the GSO dataset (Downs et al., 2022).
Ablations of multi-view generation
Our multi-view generation stage mainly consists of geometric and semantic guidance. Removing them respectively or simultaneously gives us four different combinations. As shown in Table 3 and Figure 6, the orthogonal-plane decomposition mechanism contributes the most to the geometric accuracy and consistency, bringing about visual enhancement to a great extent. The semantic guidance further increases the metric score and slightly improves visual consistency.
Table 3: Ablation studies of multi-view generation.
@@
Figure 6: Qualitative ablations of our key innovations.
Acceleration of the optimization
As mentioned in Section 3.4, we propose to use the Gaussian Divergent Significance (GDS) metric to further regularize the split and clone process. Table 5 demonstrates that this strategy has significantly reduced the optimization time while not sacrificing the reconstruction quality. Selecting the threshold carefully, our method leads to at most 15$\times$ faster convergence speed when compared with the original split and clone operation proposed in Kerbl et al. (2023).
Table 4: Quantitative results of ablating GDS metric.
[…]
w/o GDS 0.214 15min
0.01 0.168 93s
0.1 0.146 55s
0.5 0.147 78s
Table 5: Ablation studies of epipolar attention.
LPIPS $\downarrow$
w/ epipolar attn. 0.146
w/o epipolar attn. 0.231
11footnotetext: Here the reconstruction time refers to our second stage, i.e. generating 3D Gaussian representations from the multi-view images. Measured on NVIDIA V100 GPU.
Ablations of epipolar attention
To further validate the effectiveness of the proposed epipolar attention, we conduct ablation studies on it. We still adopt LPIPS to evaluate the quality of reconstruction. As presented in Table 5 and Figure 6, epipolar attention enables accurate association of features across views, fully utilizing the consistent multi-view images to synthesize high-quality 3D Gaussian Splatting.
@@ 4.4 Limitations and Future Works
While GeoGS3D shows promising results in reconstructing 3D objects from single-view images, there are still some limitations that the current framework does not entirely address. First, the number of generated views is fixed in our method. Adaptively generating different numbers of views for objects with different topological symmetries might further reduce the total reconstruction time. Additionally, our current method is restricted to single-object 3D reconstruction. It remains to be extended to complex scenes or multi-object reconstruction in the future.
@@ 5 Conclusions
In this work, we propose a two-stage model, GeoGS3D, to reconstruct 3D objects from single-view images. This method first synthesizes consistent and 3D-aware multi-view images via a diffusion model under the guidance of an orthogonal-plane decomposition mechanism. During the following 3D-GS reconstruction, epipolar attention is leveraged to communicate between relating multi-view images. A novel metric, i.e. Gaussian Divergent Significance (GDS), is proposed to accelerate optimization. Qualitative and quantitative results show that the proposed method reconstructs 3D Gaussian representations that 1) are consistent in different viewpoints, 2) are high fidelity to the reference image, and 3) display plausible creativity in the unseen areas.
Appendix A Details of Gaussian Divergent Significance (GDS)
@@
$\Upsilon_{GDS}(G(\bm{x}_{1}),G(\bm{x}_{2}))=\|\bm{\mu}_{1}-\bm{\mu}_{2}\|^{2}+tr(\bm{\Sigma}_{1}+\bm{\Sigma}_{2}-2(\bm{\Sigma}_{1}^{-1}\bm{\Sigma}_{2}\bm{\Sigma}_{1}^{-1})^{1/2}).$ (9)
The first term $\|\bm{\mu}_{1}-\bm{\mu}_{2}\|^{2}$ can be expressed as $(\bm{\mu}_{1}-\bm{\mu}_{2})^{T}(\bm{\mu}_{1}-\bm{\mu}_{2})$. Additionally, the covariance matrix can be factorized into a scaling matrix $S$ and a rotation matrix $R$: $\bm{\Sigma}=RSS^{T}R^{T}$. Since the scaling matrix is a diagonal matrix and the rotation matrix is orthogonal, we can thus simplify the computation of the second term. Given:
$\bm{\Sigma}_{1}=R_{1}S_{1}S_{1}^{T}R_{1}^{T}$ (10)
$\bm{\Sigma}_{2}=R_{2}S_{2}S_{2}^{T}R_{2}^{T},$
@@
$\quad\;tr(\bm{\Sigma}_{1}+\bm{\Sigma}_{2}-2(\bm{\Sigma}_{1}^{-1}\bm{\Sigma}_{2}\bm{\Sigma}_{1}^{-1})^{1/2})$ (11)
$=tr(\bm{\Sigma}_{1})+tr(\bm{\Sigma}_{2})-tr(2(\bm{\Sigma}_{1}^{-1}\bm{\Sigma}_{2}\bm{\Sigma}_{1}^{-1})^{1/2})$
$=tr(R_{1}S_{1}S_{1}^{T}R_{1}^{T})+tr(R_{2}S_{2}S_{2}^{T}R_{2}^{T})-tr(2(\bm{\Sigma}_{1}^{-1}\bm{\Sigma}_{2}\bm{\Sigma}_{1}^{-1})^{1/2})$
$=tr(S_{1}S_{1}^{T})+tr(S_{2}S_{2}^{T})-tr(2(\bm{\Sigma}_{1}^{-1}\bm{\Sigma}_{2}\bm{\Sigma}_{1}^{-1})^{1/2})$
@@
Appendix B Details of Epipolar Line